# Modeling Binary Lenses and Sources with the BAGLE Python Package

### v1

@@ Appendix C $t_{0}$ transformation
$S^{\prime}-S=dcos\phi\hat{\mu}_{rel}$ (C3)
Plugging this into Eq. C1, we find
$dcos\phi\hat{\mu}_{rel}$ $=$ $\frac{\boldsymbol{\mu_{rel}}}{\theta_{E}}(t_{0}^{\prime}-t_{0})$ (C4)
$dcos\phi\hat{\mu}_{rel}$ $=$ $\frac{1}{t_{E}}(t_{0}^{\prime}-t_{0})\hat{\mu}_{rel}$ (C5)
$\boxed{t_{0}^{\prime}=t_{0}+t_{E}dcos\phi}$ (C6)
[…]
C.1 Standard Coordinate Transforms
As described in Section B.1, for a geometric midpoint $\leftrightarrow$ primary center transformation Eq. C6 becomes
$t_{prim,0}=t_{geom,0}+t_{E}\frac{a}{2\theta_{E}}cos\phi$ (C7)
and for a geometric midpoint $\leftrightarrow$ center of mass transformation Eq. C6 becomes
$t_{prim,0}=t_{geom,0}+t_{E}sq^{\prime}cos\phi$ (C8)

### v3

@@ Appendix D $t_{0}$ transformation
$S^{\prime}-S=\hat{\mu}_{rel}d\cos\phi$ (D3)
Plugging this into Eq. D1, we find
$\hat{\mu}_{rel}d\cos\phi$ $=$ $\frac{\boldsymbol{\mu_{rel}}}{\theta_{E}}(t_{0}^{\prime}-t_{0})$ (D4)
$\hat{\mu}_{rel}d\cos\phi$ $=$ $\frac{1}{t_{E}}(t_{0}^{\prime}-t_{0})\hat{\mu}_{rel}$ (D5)
$\boxed{t_{0}^{\prime}=t_{0}+t_{E}d\cos\phi}$ (D6)
Figure 19: Similar to Figure 18 but with the angle, $\phi$ between $\boldsymbol{\mu_{rel}}$ and the binary axis ($\boldsymbol{s}$) marked. $\phi<90^{\circ}$ chosen for ease of visualization. The projection of $d\hat{s}$ onto $\boldsymbol{\mu_{rel}}$ is $\hat{\mu}_{rel}d\cos\phi$.
D.1 Standard Coordinate Transforms
As described in Section C.1, for a geometric midpoint $\leftrightarrow$ primary center transformation Eq. D6 becomes
$t_{prim,0}=t_{geom,0}+t_{E}\frac{a}{2\theta_{E}}\cos\phi$ (D7)
and for a geometric midpoint $\leftrightarrow$ center of mass transformation Eq. D6 becomes
$t_{prim,0}=t_{geom,0}+t_{E}sq^{\prime}\cos\phi$ (D8)